prhit clearly): 14) (10 pts) In humans is the number of babies born in Hawaii and most other mammals, the sex ratio is expected to be 1:1 male:female. Below Observed Male per 100 live births 90 during 1998. Observed Female per 100 live births TOTAL births 86 17,566 a. List the variables and data types b. Write out an appropriate null hypothesis? c. Name and BRIEFLY justify the statistical test you recommend for test d. What are the degrees of freedom for a test of the null hypothesis? e. The test statistic was 0.5882. Do you provisionally accept or reject the null hypothesis? Explain your choice
Answer #1

Answer:

A) no of male(n1) = no of female(n2) = 17566/2 = 8783

p1 = 0.9

p2 = 0.86

B) H0: p1 = p2

H1: p1 \neq p2

C) It is a hypothesis test of proportion and we will take Z-test.

D) Degrees of freedom (DF) = n - 1

E) At alpha = 0.05, the critical value is 1.96

As the test statistic value 0.5882 is greater than the critical value so the null hypothesis is rejected.